# A Monte Carlo Tree Search approach to QAOA: finding a needle in the haystack

### v2

@@ V Conclusions and outlook
In this paper, we analyzed the performance of MCTS as a gradient-free optimizer for QAOA variational parameters. Despite its success in designing optimal schedules for quantum annealing [36], a vanilla MCTS fails when faced with the complex energy landscape of QAOA. This complexity emerges as fractal-like decision trees, characterized by the close proximity of low and high-energy leaf nodes, where it is impossible to discriminate between good and bad branches in polynomial time reliably. We overcame these difficulties by restricting the search space of QAOA parameters based on the knowledge of optimal parameters obtained for shallower circuits and leveraging the well-known presence of regular schedules in QAOA [9, 19, 28, 46, 47, 25]. By testing it on 3-SAT and MaxCut instances, we showed that iterative SSR-MCTS provides a systematic way to optimize QAOA parameters within the MCTS framework. Its performance is comparable with other heuristic optimization methods while offering at the same time exceptional resilience against noise, suggesting it is perfectly suited for NISQ applications.
@@
Moreover, the flexibility of MCTS suggests further possible development. First, further modifications to the tree policy can be envisioned to improve the performance and tailor it to specific VQA applications. For instance, the algorithm would likely benefit from changes to the selection criterion itself [49], such as using a different kind of policy entirely [75, 76] or modifying the UCB1 to favor the exploration of nodes with a large reward variance [50, 77]. MCTS can also be efficiently integrated with neural networks [39, 36] to speed up the nodes’ score evaluations and improve the transferability between different problem instances or sizes. This will likely make the algorithm more robust at the price of a greater computational cost in the training phase. Finally, we expect MCTS to be paired well with many modifications of QAOA as well as other VQAs. Being problem agnostic and relying only on a well-defined reward (cost) function, which is a key ingredient of VQAs, iterative SSR-MCTS represents a reliable workhorse to add to the toolkit of hybrid quantum-classical algorithms.
Optimizing VQA parameters, in general, is known to be a hard task as features like barren plateaus [34] and proliferation of local minima [26] hinder the exploration of the energy landscape. Warm initialization of the variational parameters has often been shown to cure these problems [78, 79, 80] and, essentially, iterative SSR-MCTS falls on the same line. Despite some common ground, standard variational optimization and MCTS tackle the problem from very different perspectives: the former sees a function defined on a large manifold and looks for its local or global minima, and the latter approaches QAOA as a sequential decision-making process. Hence, an interesting perspective is whether it is possible to connect the “fractalization” of the decision tree in MCTS with the structure of the energy landscape itself. In other words, is the self-similar pattern of the leaf nodes (see Appendix  B) correlated to the complexity (roughness or flatness) of the energy landscape [53]? Indeed, one might argue that the complexity of the optimization task is related more to the QAOA algorithm itself and less to the way we optimize its parameters; a hard problem remains hard irrespective of the strategy used to tackle it. On the other hand, it has been shown [28] that a reinforcement learning approach can identify regular patterns in QAOA parameters that are otherwise hard to find with more standard warm initializations of local optimizations. Therefore, it is possible that also MCTS, being a crude reinforcement learning algorithm, might show an advantage over direct minimization in the energy landscape for certain problems where standard optimizers struggle the most. For instance, Bayesian optimization [81] represents a compelling alternative that has recently gained momentum as a powerful and robust algorithm for the classical loops of VQAs [82, 83]. A detailed comparison of their respective strengths and limitations will contribute towards the realization of a comprehensive toolkit for hybrid quantum-classical algorithms, where the most appropriate optimizer can be chosen depending on the task at hand. Addressing these issues will advance our understanding of the mutual enhancement between quantum computing and artificial intelligence, and foster the development of quantum technologies.
Acknowledgments
@@ Appendix B Why standard MCTS struggles: structure of the decision tree with and without SSR
Visually, the two cases are vastly different. The unrestricted case shows a very complex and unpredictable distribution of leaf nodes, where good strategies are likely hard to find. The best nodes are hidden in regions where very badly scoring ones also exist, and the averaging of the UCB1 strategy makes standard MCTS blind to these features in favor of paths that lead to mediocre leafs more consistently. The improvements due to SSR can be argued qualitatively already from the structure of the energies in Fig. 10. There is one key observation: to restrict this $P=2$ search space, we make use of the optimal parameters $(\tilde{\gamma}_{1}^{*},\tilde{\beta}_{1}^{*})$ for the same instance at $P=1$. As such, $\tilde{\gamma}_{1}^{*}$ sets an upper bound on $\gamma_{1}$, which means that the entire right part of the leafs (those with $\gamma_{1}>\tilde{\gamma}^{*}_{1}$) in Fig. 10 will be discarded. Indeed, we see that these turn out to all be undesirable leaf nodes. Furthermore, $\tilde{\beta}_{1}^{*}$ sets a lower bound for $\beta_{1}$, which in turn means that the left part within each first-move sector is filtered out, and, again, we see that it is the parts with high $\beta_{1}$ that house the best minima. This qualitative argument can be extended all the way to the choice of the last parameter.
@@ Appendix C MCTS variants in the midst of noise
Figure 12: Comparison of the performance of different MCTS variants with varying noise strength. Average energy obtained with the standard max child (blue circles), robust child (orange diamonds) and single-player variant (green squares), where the child leading to the highest obtained reward is chosen. The crosses of each color represent the best game for each noise setting for the corresponding variant. We also show the lowest energy obtained by COBYLA (black crosses) among 20 runs for each noise strength (only up to 0.4). The SSR-MCTS results are averaged over 15 games for each noise configuration. The results correspond to an example 3-SAT instance with $n=7$. The horizontal red and green lines indicate the optimal energy given by the $P=1$ and $P=2$ optimal angles, respectively.
In the main text, we utilized standard SSR-MCTS (i.e., without the single-player modifications) to optimize noisy cost functions (Sec. III.4). In this appendix, we briefly look at how other versions of our algorithm fare within noisy landscapes. Apart from the standard MCTS algorithm (the one which makes the final decision by choosing the child with maximum $w_{i}/n_{i}$, also known as max child), we also test the standard algorithm with another widespread final move selection criterion known as robust child [88, 41, 89], which chooses the child with the highest visit count $n_{i}$, as well as the single-player variant introduced in Sec. II.2. We also compare the different MCTS approaches with COBYLA [69], a gradient-free optimizer widely used in VQAs and as a standard baseline comparison to other optimizers [90, 83, 91], despite not being particularly optimized for noisy landscapes [83, 92].
@@ Appendix D Resource cost analysis
As an anytime algorithm [41], the number of times the cost function is evaluated $n_{\rm fev}$ can be very easily controlled with MCTS. The number of calls matches the number of MCTS cycles performed during the game, times the number of shots needed for the desired accuracy, and they can be stopped at any point to make the algorithm suggest a strategy. The minimum number of cycles that we can run without leaving out any options are $2Pb$ (the depth of the tree times the branching factor), where the selected move at each turn will be the child node which gave the highest score after running one random simulation from each of the $b$ children. Evidently, the results from such MCTS runs are bound to be very inaccurate. We also know that, as the number of cycles grows to infinity, the algorithm eventually finds the optimal leaf node [41]. In general, the training of machine-learning and AI methods is resource expensive and we expect MCTS to be no exception. In the main text, we used a heuristically chosen scaling $n_{\rm fev}=1000+800(2P-1)$. Here we elucidate the performance dependence on $n_{\text{fev}}$, which can be used to balance between the algorithm’s cost and its accuracy. Importantly, the $n_{\text{fev}}$ cycles do not need to be evenly distributed among the $2P$ choices that constitute the final strategy. Indeed, we find that spending more time at the beginning of the process is beneficial, as the first moves are critical for designing a successful schedule. Instead, optimizing the parameters “deeper” in the Anstaz requires less exploration of the remaining part of the tree.
@@
Figure 13 shows the performance of iterative SSR-MCTS with varying number of total MCTS cycles and system sizes. The results in Fig. 13(a) are obtained with $\delta=0$ on a graph with $n=6$ vertices (shown below the main plot), while the softening is $\delta=0.2$ in Fig. 13(b), where we run the algorithm with $P=7$ layers on a graph with $n=10$ (also shown below the plot). The branching factor is set to $b=30$ in both cases. The sizes of the trees in the two cases thus amount to $8.1\times 10^{5}$ and $2.2\times 10^{10}$ leaf nodes, respectively, out of which the algorithm samples $n_{\text{fev}}$ (not necessarily unique) nodes. The two panels display an improving performance as $n_{\text{fev}}$ increases, both in the average and in the best-found approximation ratio. However, the range in which $r$ varies is rather limited, below $2\%$, and it would be significantly overshadowed by the presence of noise. Hence, a few hundred cycles are enough to reach good approximations of the target state, as the search-space restriction already limits the variational Anstaz in a region close to the global minimum. In particular, notice that in Fig. 13(b) SSR-MCTS already reaches $r\geq 0.95$ for the minimum number of cycles considered $n_{\text{fev}}=420$. A comparison with other optimization strategies [82] shows that iterative SSR-MCTS is competitive against other state-of-the-art algorithms.
@@
A resource-savvy implementation will invest more time in the optimization of low-depth ansätze and progressively reduce the number of cycles per layer as $P$ increases up to the desired accuracy or until the energy no longer decreases. If a given QAOA setup is particularly limited on function evaluations, the branching factor could also be reduced to make the algorithm converge with fewer calls at the cost of the reachability of the SSR-MCTS algorithm. These coarser results could also be used as good initial parameters for gradient-based runs (see App. A).

### v3

@@ V Conclusions and outlook
In this paper, we analyzed the performance of MCTS as a gradient-free optimizer for QAOA variational parameters. Despite its success in designing optimal schedules for quantum annealing [39], vanilla MCTS fails when faced with the complex energy landscape of QAOA. This complexity emerges as fractal-like decision trees, characterized by the close proximity of low- and high-energy leaf nodes, where it is impossible to discriminate between good and bad branches in polynomial time reliably. We overcame these difficulties by restricting the search space of QAOA parameters based on the knowledge of optimal parameters obtained for shallower circuits and leveraging the well-known presence of regular schedules in QAOA [11, 21, 30, 49, 50, 27]. By testing it on 3-SAT and MaxCut instances, we showed that iterative SSR-MCTS provides a systematic way to optimize QAOA parameters within the MCTS framework. Preliminary numerical results indicate its performance is comparable with other heuristic optimization methods while offering at the same time great resilience against noise, suggesting it effectively accommodates both the potentiality and limitations of NISQ applications.
@@
Moreover, the flexibility of MCTS suggests further possible development. First, further modifications to the tree policy can be envisioned to improve the performance and tailor it to specific VQA applications. For instance, the algorithm would likely benefit from changes to the selection criterion itself [52], such as using a different kind of policy entirely [81, 82] or modifying the UCT to favor the exploration of nodes with a large reward variance [53, 83]. MCTS can also be efficiently integrated with neural networks [42, 39] to speed up the nodes’ score evaluations and improve the transferability between different problem instances or sizes. This will likely make the algorithm more robust at the price of a greater computational cost in the training phase. Finally, we expect MCTS to be paired well with many modifications of QAOA as well as other VQAs. Being problem agnostic and relying only on a well-defined reward (cost) function, which is a key ingredient of VQAs, iterative SSR-MCTS represents a reliable workhorse to add to the toolkit of hybrid quantum-classical algorithms.
Optimizing VQA parameters is generally a challenging task due to features such as barren plateaus [36, 37] and the proliferation of local minima [28], which hinder efficient exploration of the energy landscape. Warm initialization of variational parameters has been shown to mitigate these difficulties [84, 85, 86], and iterative SSR-MCTS essentially follows a similar philosophy. Despite this common ground, standard variational optimization and MCTS approach the problem from fundamentally different perspectives: The former treats QAOA as a function optimization problem over a high-dimensional manifold, while the latter frames it as a sequential decision-making process.
This distinction motivates further studies to investigate how the structural properties of the decision tree in MCTS relate to the geometry of the energy landscape itself. In particular, could the observed “fractalization” of the decision tree be indicative of the roughness or flatness of the energy landscape [56]? Understanding such connections could offer deeper insights into the intrinsic difficulty of variational optimization, as well as reveal new ways to exploit decision-tree structures to navigate complex landscapes more effectively. Indeed, one might argue that the hardness of QAOA optimization is an inherent property of the algorithm rather than a consequence of the optimization method used. However, reinforcement learning approaches have been shown to uncover regular patterns in QAOA parameters that remain elusive to traditional warm-started optimizations [30]. Given that MCTS can be viewed as a crude form of reinforcement learning, it is possible that it could outperform direct energy landscape minimization in certain cases where conventional optimizers struggle the most.
Furthermore, alternative strategies such as Bayesian optimization [87] and surrogate methods [88, 89] have recently gained traction as powerful and robust tools for the classical loops of VQAs [90, 91, 92]. A detailed comparison of their respective strengths and limitations, as well as an investigation of possible fruitful combinations of these methods, would be instrumental in developing a more comprehensive toolkit for hybrid quantum-classical algorithms, enabling the selection of the most effective optimization method for a given problem.
Another crucial aspect in the development of quantum algorithms for NISQ devices is their resilience against errors and decoherence. These, indeed, typically degrade the performance of VQAs when the circuit depth increases, even if the optimization loop is entirely done on classical machines [93]. For more realistic applications, it is therefore important to understand whether iterative SSR-MCTS can cope with hardware noise and still suggest meaningful strategies.
Addressing these issues will advance our understanding of the mutual enhancement between quantum computing and artificial intelligence, and foster the development of quantum technologies.
Data availability statement
The data that support the findings of this study are openly available at [94].
Acknowledgements.
@@ Appendix B Why standard MCTS struggles: structure of the decision tree with and without SSR
Visually, the two cases are vastly different. The unrestricted case shows a very complex and unpredictable distribution of leaf nodes, where good strategies are likely hard to find. The best nodes are hidden in regions where very badly scoring ones also exist, and the averaging of the UCT strategy makes standard MCTS blind to these features in favor of paths that lead to mediocre leafs more consistently. The improvements due to SSR can be argued qualitatively already from the structure of the energies in Fig. 9. There is one key observation: to restrict this $P=2$ search space, we make use of the optimal parameters $(\tilde{\gamma}_{1}^{*},\tilde{\beta}_{1}^{*})$ for the same instance at $P=1$. As such, $\tilde{\gamma}_{1}^{*}$ sets an upper bound on $\gamma_{1}$, which means that the entire right part of the leafs (those with $\gamma_{1}>\tilde{\gamma}^{*}_{1}$) in Fig. 9 will be discarded. Indeed, we see that these turn out to all be undesirable leaf nodes. Furthermore, $\tilde{\beta}_{1}^{*}$ sets a lower bound for $\beta_{1}$, which in turn, means that the left part within each first-move sector is filtered out, and, again, we see that it is the parts with high $\beta_{1}$ that house the best minima. This qualitative argument can be extended all the way to the choice of the last parameter.
@@ Appendix C MCTS variants in the midst of noise
Figure 11: Comparison of the performance of different MCTS variants with varying noise strength. Average energy obtained with the standard max child (blue circles), robust child (orange diamonds) and single-player variant (green squares), where the child leading to the highest obtained reward is chosen. The crosses of each color represent the best game for each noise setting for the corresponding variant. We also show the lowest energy obtained by COBYLA (black crosses) among 20 runs for each noise strength (shown only up to 0.4). The SSR-MCTS results are averaged over 15 games for each noise configuration. The results correspond to an example 3-SAT instance with $n=7$. The horizontal red and green lines indicate the optimal energy given by the $P=1$ and $P=2$ optimal angles, respectively.
In the main text, we utilized standard SSR-MCTS (i.e., without the single-player modifications) to optimize noisy cost functions (Sec. III.4). In this appendix, we briefly look at how other versions of our algorithm fare within noisy landscapes. Apart from the standard MCTS algorithm (the one which makes the final decision by choosing the child with maximum $w_{i}/n_{i}$, also known as max child), we also test the standard algorithm with another widespread final move selection criterion known as robust child [99, 44, 100], which chooses the child with the highest visit count $n_{i}$, as well as the single-player variant introduced in Sec. II.2. We also compare the different MCTS approaches with COBYLA [75], a gradient-free optimizer widely used in VQAs and as a standard baseline comparison to other optimizers [101, 92, 102], despite not being particularly optimized for noisy landscapes [92, 103]. We do not provide an explicit comparison with BFGS, a standard gradient-based benchmark algorithm for noiseless optimization, as we found it unable to converge even at the smallest noise setting that we considered.
@@ Appendix D Resource cost analysis
As an anytime algorithm [44], the number of times the cost function is evaluated $n_{\rm fev}$ can be very easily controlled with MCTS. The number of calls matches the number of MCTS cycles performed during the game, times the number of shots needed for the desired accuracy, and they can be stopped at any point to make the algorithm suggest a strategy. The minimum number of cycles that we can run without leaving out any options are $2Pb$ (the depth of the tree times the branching factor), where the selected move at each turn will be the child node which gave the highest score after running one random simulation from each of the $b$ children. Evidently, the results from such MCTS runs are bound to be very inaccurate. We also know that, as the number of cycles grows to infinity, the algorithm eventually finds the optimal leaf node [44]. In general, the training of machine-learning and AI methods is resource expensive and we expect MCTS to be no exception. In the main text, we used a heuristically chosen scaling $n_{\rm fev}=1000+800(2P-1)$. Here we elucidate the performance dependence on $n_{\text{fev}}$, which can be used to balance between the algorithm’s cost and its accuracy. Importantly, the $n_{\text{fev}}$ cycles do not need to be evenly distributed among the $2P$ choices that constitute the final strategy. Indeed, we find that spending more time at the beginning of the process is beneficial, as the first moves are critical for designing a successful schedule. Instead, optimizing the parameters “deeper” in the ansatz requires less exploration of the remaining part of the tree.
@@
Figure 12 shows the performance of iterative SSR-MCTS with varying number of total MCTS cycles and system sizes. The results in Fig. 12(a) are obtained with $\delta=0$ on a graph with $n=6$ vertices (shown below the main plot), while the softening is $\delta=0.02$ in Fig. 12(b), where we run the algorithm with $P=7$ layers on a graph with $n=10$ (also shown below the plot). The branching factor is set to $b=30$ in both cases. The sizes of the trees in the two cases thus amount to $8.1\times 10^{5}$ and $2.2\times 10^{10}$ leaf nodes, respectively, out of which the algorithm samples only $n_{\text{fev}}$ (not necessarily unique) nodes. The two panels display an improving performance as $n_{\text{fev}}$ increases, both in the average and in the best-found approximation ratio. However, the range in which $r$ varies is rather limited, below $2\%$, and it would be significantly overshadowed by the presence of noise. Hence, a few hundred cycles are enough to reach good approximations of the target state, as the search-space restriction already limits the variational ansatz in a region close to the global minimum. In particular, notice that in Fig. 12(b) SSR-MCTS already reaches $r\geq 0.95$ for the minimum number of cycles considered $n_{\text{fev}}=420$. Our results suggest that iterative SSR-MCTS is competitive against other state-of-the-art algorithms, such as the Bayesian Optimization in ref. [90]. The performance can be compared on the same Max-Cut instance ($n=10$ and $P=7)$, where SSR-MCTS obtains a better approximation ratio. A comparison of the respective resource costs, however, is less straightforward as our algorithm leverages the knowledge of optimal schedules on shallower circuits and, therefore, it depends on the details of the implementation of the iterative procedure.
@@
A resource-savvy implementation will invest more time in the optimization of low-depth ansätze and progressively reduce the number of cycles per layer as $P$ increases up to the desired accuracy or until the energy no longer decreases. In cases where the number of function evaluations is an important limiting factor of a particular experimental implementation of QAOA, the branching factor could also be reduced to make the algorithm converge with fewer calls at the cost of the reachability of the SSR-MCTS algorithm. These coarser results could also be used as good initial parameters for gradient-based runs (see Appendix A). Additionally, integrating MCTS with neural networks, as done in Ref. [39] for annealing schedules, will likely greatly enhance our strategy. Indeed, the neural networks allow for adapting existing trained MCTS agents on new problem instances or circuit depths without having to learn the decision tree structure from scratch.
A separate discussion concerns the number of shots needed for any function evaluation. To estimate the reward with an accuracy $\epsilon$, typically one needs $N_{\rm shots}=O(1/\epsilon^{2})$ single-shot measurements. Here, SSR-MCTS has two nice features that contribute towards reducing this overhead. First, the robustness against measurement noise suggests that rough estimates of the reward are sufficient for obtaining good approximations in the ground states, in particular in the early training stage of the MCTS agent. Second, The prefactor in the proportionality relation between $N_{\rm shots}$ and $\epsilon^{-2}$ will likely decrease as $P$ increases. Indeed, SSR already severely restricts the energy window spanned by the QAOA ansatz, as clearly shown by Fig. 3. Their energy variance is reduced accordingly. Finally, one can train the MCTS algorithm directly with single-shot measurements and update the value of the tree nodes using the energy of the sampled classical configuration. This might greatly reduce the resource requirements, but further investigations are needed to gauge its efficacy and possible advantage over the approach presented here.